OpenStudy (anonymous):

Write the partial fraction decomposition of the rational expression. Check your result algebraically. x2 + 6x + 3 x3 + x

OpenStudy (anonymous):

edit (x^2+6x+3)/(x^3+x)

OpenStudy (anonymous):

i got 30i+75j+0k

OpenStudy (anonymous):

1/sqareroot of 6525 <30,75,0>

OpenStudy (anonymous):

\[\frac{x^2+6x+3}{x(x^2+1)}=\frac{A}{x}+\frac{Bx+C}{x^2+1}\]

OpenStudy (anonymous):

you must be doing a different problem

OpenStudy (anonymous):

wait you are right

OpenStudy (anonymous):

but thanks for that one too